# The galaxy-IGM connection in thesan: observability and information content of the galaxy-Lyman- cross-correlation at z6

### v1

@@ 3 Impact of observational limitations
The leading interpretation of the GaLCC paints a picture where the observed signal arises from two opposite effects, each dominating at different scales around galaxies. The enhanced ionizing photon density provided by (large) galaxies boosts the transmission of Ly in their surroundings. This transverse proximity effect can be detected statistically by looking at the excess Ly flux around galaxies. At the same time, the overdensity where galaxies reside boosts the HāII recombination rate, suppressing the Ly flux. This is the dominant effect at the smallest scales, while the aforementioned radiation proximity effect takes over at intermediate ones. In Fig.Ā 2 we visually show this by reporting in the top panel the GaLCC computed from thesan-1 at and using galaxies with stellar mass (black solid line). We mark the radii corresponding to , , , and Mpc with light grey vertical solid lines. In the middle and bottom panels, we show the stacked ionizing photon density and gas density around the same galaxies used in the computation of the GaLCC. In order to compute these quantities, we extract gas properties in a cubic region with side Mpc and centered on each galaxy in the sample. We then stack these regions and finally project the resulting cube along one axis for visualisation purposes. In the middle and bottom panel, we mark with dashed circles the same radii as above. It can clearly be seen that the central regions (approximately within - Mpc) are dominated by high densities, which suppress the Ly transmission thanks to the boosted recombination rate suppressing the ionised fraction. At intermediate scales () the density has already fallen to the background level, but the ionizing photons density is still largely above the one observed farther away from the center. The latter enhances the Ly flux at such scales, resulting in a broad peak in the GaLCC.
@@ 3.1 How many spectra are necessary to measure the GaLCC?
The first question we address is the following: How many spectra do we need to obtain statistically-sound results? The inhomogeneous nature of reionization implies that the gas ionization state and temperature in different regions of the Universe vary significantly until (Bolton etĀ al., 2017). Additionally, the highly non-linear process of galaxy formation, the consequent energy injection (āfeedbackā) into the surrounding circum-galactic medium (CGM) and IGM, as well as the different cosmic environments hosting galaxies of similar mass all can strongly influence the Ly forest signal. Therefore, individual lines of sight are dominated by such āstructure formation noiseā. Current observational efforts are still limited to a handful of sightlines, that however might be not sufficient to extract the intrinsic signal.
@@
In order to test the reliability of current constraints and to guide future observations, we show in Fig.Ā 3 the GaLCC signal extracted from our simulations for different values of the total length of the Ly forest spectra (). In practice, we employ up to 300 spectra of equal length , but decide to report the results using the total spectral length in order to ease the comparison with observations. In each panel, the thin colored lines are the values obtained using different sets of sightlines with total spectral length equal to the value reported in the bottom right part of the panel. For comparison, we show using a thick black line our āgolden standardā, i.e.Ā the value obtained using all 300 spectra available. It appears immediately clear that this signal is noise-dominated in the top panels (corresponding to ), while the recovered signal becomes close to the intrinsic one (that we assume is traced by our black thick line) for .
@@
We compare the prediction from our simulation with available observations, namely Meyer etĀ al. (2019, cGpc), Meyer etĀ al. (2020, cGpc), and Kashino etĀ al. (2023, cMpc). Observations from Meyer etĀ al. (2019, 2020) span a large enough path length to extract the true reionization signal. It should be noted, however, that to reach such large , they compound together large redshift intervals, potentially encountering lightcone effects due to the rapid evolution of the ionization field towards the end of the EoR (as discussed in Sec.Ā 3.6). In the case of Kashino etĀ al. (2023), reporting the results from a single sightline (out of six expected) from the EIGER survey, we predict that their results are completely dominated by the āstructure formation noiseā, and therefore cannot be quantitatively nor qualitatively trusted.
The aforementioned results put the findings of Kashino etĀ al. (2023) in a different light. In fact, while they find a qualitative similarity between their measured and the one reported by the thesan simulations at redshift , such agreement is likely entirely driven by noise. However, with the completion of the EIGER survey (six quasar fields), we predict that the noise level will be reduced enough to enable qualitative, if not quantitative, statements.
@@
It should be noted that quantitative results depend on the galaxies selected when computing the cross-correlation (or, similarly, detected by a survey). In particular, in Fig.Ā 3 we have selected all galaxies with stellar mass . We provide a visual impression of the impact of such choice in Fig.Ā 4 and 5, where we have selected all galaxies with stellar mass and , respectively. Comparing these figures it appears clear that larger threshold masses increase the noise in the recovered signal because of the diminishing number of galaxies selected with higher mass thresholds, therefore pushing the requirements to longer in order to maintain under control the noise. We have checked that this conclusion is negligibly affected by the choice of binning.
@@ 3.2 What field of view do we need?
Many observational surveys have limited field of view around the sightline. This can negatively impact their ability to study the GaLCC, since it requires us to probe distances up to cMpc in order to capture the extent of the flux enhancement. However, it is important to realise that such distance is not the transverse distance to the sightline, but rather the 3D distance between a pixel in the spectrum and a galaxy. Therefore, even when observations probe only a smaller transverse distance (), they can still probe 3D separations , at the price of a smaller statistic. Therefore, here we answer the question: What is the impact of on the cross-correlation signal?
In order to address this question, we show in Fig.Ā 6 the cross-correlation obtained by imposing different values for the maximum galaxy-sightline (transverse) distance (solid colored lines). For comparison, we show the predicted cross-correlation with no maximum limit (black dashed line). It appears clear that a transverse length as small as is sufficient to capture the cross-correlation signal without adding any significant noise. However, limiting severely decreases the number of configurations (i.e.Ā galaxy-pixel pairs) at distances , therefore increasing the statistical noise. Nevertheless, our results are very encouraging for future surveys aiming at detecting this cross-correlation, as even with limited fields of view around the quasar sightline, the signal can be faithfully recovered. In fact, while both probing larger and observing more sightlines increase the number of galaxy-pixel pairs, the latter allow us to probe a larger variety of cosmic environments, and it should therefore be favoured.
@@ 3.3 How accurate should the spectra be?
We begin by exploring the impact of noise, assumed to be Gaussian and uncorrelated between pixels. The noise level is characterised by the signal-to-noise ratio (S/N) of the continuum-normalized spectrum (i.e.Ā the signal is the normalized flux). We do not impose the resulting flux in each pixel to be positive (but show in AppendixĀ B the implication of doing so). For each value of the S/N, we produce 50 different realisations of the Gaussian noise and show in Fig.Ā 7 the range of values that the cross-correlation can assume (shaded regions with color reflecting the S/N). The figure shows that the cross-correlation signal is well recovered for S/N as low as 1% of the normalised flux. However, this results emerges only when showing the full envelope of possible values, while for a single noise realisation the true signal is almost-always completely lost in the noise, as discussed already.
@@
We note that in this specific case, the choice of focusing on the signal at has some impact. In fact, this redshift marks a sweet spot in the competition between a larger intrinsic signal (achieved by moving to earlier times Garaldi etĀ al., 2022) and more numerous galaxies to beat down the statistical noise (found at later times due to structure growth). Therefore, the noise requirements will become more stringent at different times, either because of the growing impact of statistical noise or because the intrinsic signal weakens (i.e.Ā the cross-correlation flattens on to the line).
@@
The next property of the spectra that we investigate is the spectral resolution , where is the largest velocity difference that can be told apart and is the speed of light. For each of the investigated values of we have rebinned the spectra using a boxcar filter to achieve the desired resolution (the original synthetic spectra have resolution km/s). Fig.Ā 8 shows the resulting cross-correlation for the different values, and clearly shows that a spectral resolution of is sufficient to capture the salient features of the cross-correlation at radii . A somewhat more stringent constraints is found at smaller separations (), where a spectral resolution of is required in order to recover the intrinsic cross-correlation. The reason is simply that at such resolution a pixel corresponds to a physical length of , and therefore smaller scales cannot be captured by the cross-correlation.
@@
The loose resolution requirements are very promising, as it renders much easier to increase the number of observed spectra. Unfortunately, the impact of resolution and noise are not independent. In fact, the resilience to noise comes from the averaging process performed while computing the cross-correlation, which averages out the impact of pixel-level noise. However, this only works as long as the regions of enhanced and suppressed transmitted flux in the spectrum contain a sufficient number of pixels. Therefore, when the spectral resolution decreases the resilience to noise worsen. To demonstrate this and to ease the task of striking a balance between the spectral resolution and S/N when attempting to recover the intrinsic signal, we show in Fig.Ā 9 the impact of noise spectra of different resolution. In practice, we explore the S/N in Fig.Ā 7 (colored bands) and spectral resolutions in Fig.Ā 8 (panels from top to bottom, is indicated in the bottom right of each one of them). As anticipated, lower are more sensitive to noise. However, even for spectral resolutions as low as (second panel from the top), a S/N of 20 (dark blue band) is sufficient to recover the general GaLCC shape, while for (third panel from the top), S/N=10 is sufficient. We remind the reader that the colored bands show the outer envelope of the GaLCC computed for different noise realizations, therefore providing a worst-case scenario. In reality, most of the observed curves will be significantly closer to the noiseless one.
@@
Our results show that even with significant noise it is possible to recover the salient features of the cross-correlation signal, shining a promising light on the possibility of significantly enlarging the number of los usable for this measurement. Clearly, a major issue in using sightlines for this purpose is the need to identify galaxies around them. Overall, these results, combined with the one provided in Sec.Ā 3.1, outline the observational requirements for the reliable characterization of the GaLCC. In Sec.Ā 5 we will discuss in more details which surveys (available or forthcoming) can be used for this measurement.
@@ 3.4 How to select galaxies?
Observations of the GaLCC employ different techniques to identify galaxies, ranging from CIV absorption in the same spectrum to OāIII emission around the sightline. Therefore, it is unclear how to properly compare such observations among themselves and with simulations. For this reason, here we have decided to remain agnostic and select galaxies based on their stellar mass. However, it is keyto determine whether these different selection techniques are biasing the inferred GaLCC. Additionally, it is foreseeable that in the coming years more and more surveys will identify galaxies through their emission lines. The OāIII emission line is particularly promising, and is in fact used by a number of ongoing JWST surveys. For this reason, we study here the effect of selecting galaxies based on their star formation rate (SFR), stellar mass () and OāIII flux. For the latter, we employ the synthetic galaxy SEDs described in Kannan etĀ al. (2022b); Garaldi etĀ al. (2024). Thanks to these, we can restrict the galaxy selection to the brightest OāIII emitters in the simulation box, defined here based on their equivalent width (EWOIII). While the limited volume of our simulations hinders a full one-to-one comparison with e.g.Ā the EIGER survey, it still provides precious information on the biases carried by such selection. In Fig.Ā 10 we show the cross-correlation computed selecting the top 0.1%, 1%, 20% and 50% (panels from top to bottom) galaxies when ranked in OāIII flux, SFR, stellar mass and gas metallicity (as a simple proxy for the strength of their CāIV absorption signature in the QSO spectrum). Notice that the galaxy SED were computed only for a subset of āwell-resolvedā galaxies (as defined in Kannan etĀ al., 2022b), resulting in approximately 10000 galaxies at . Therefore, to enable a proper comparison, we have restricted all selection to the same subset of galaxies, despite the SFR, stellar mass and gas metallicity being available for all simulated galaxies. The different panels show a consistent picture, namely that the selection performed does not affect the resulting cross-correlation. More selective criteria result in a more noisy signals (because of the smaller number of objects resulting in larger noise) and higher peaks (because the sources are more biased and more luminous, as discussed in Sec.Ā 4.1 Garaldi etĀ al., 2022), but there is no discernible difference between selecting galaxies based on their stellar mass, star formation rate or OāIII flux.
@@
The insensitivity of the GaLCC to the source selection stems from two conspiring reasons. First, in the thesan simulation most galaxies lie on the star-formation main sequence and on the mass-metallicity relation (Garaldi etĀ al., 2024), with only very few deviating from it (Shen etĀ al., 2024). Therefore, the most star forming galaxies are also the most massive ones (in terms of stellar component) and the most metal-rich ones. Second, the large volume covered by the simulations does not allow to resolve the ISM of the simulated galaxies. Therefore, the line emission has to be āpaintedā on the unresolved ISM using approximate methods (Kannan etĀ al., 2022b). This entails that, for the OāIII line, there is a built-in correlation between its flux and the star formation of the galaxy. While the complex radiation transport and dust obscuration entering in the SED production can break this correlation, it is the case that the OāIII flux and the galaxy SFR are tightly related in thesan (Kannan etĀ al., 2022b). These two effects conspire to essentially yield the same galaxy sample, regardless of the selection method. In fact, the overlap between the samples of galaxies resulting from the three selection methods discussed is larger than 85% in all but one case. We caution, however, that this result might change if we were able to resolve the ISM of galaxies in a large-enough volume to compute the GaLCC. Unfortunately this is currently beyond the reach of current simulations.
@@
The results presented above strengthen the approach taken in this paper and in previous theoretical analysis, i.e.Ā selecting galaxies based on their stellar mass. It also provides precious information for future observational studies.
3.5 Impact of gaps in the galaxy catalog

### v2

@@ 3.1 How many spectra are necessary to measure the GaLCC?
The first question we address is the following: How many spectra do we need to obtain statistically-sound results? The inhomogeneous nature of reionization implies that the gas ionization state and temperature in different regions of the Universe vary significantly until (Bolton et al., 2017). Additionally, the highly non-linear process of galaxy formation, the consequent energy injection (‘feedback’) into the surrounding circum-galactic medium (CGM) and IGM, as well as the different cosmic environments hosting galaxies of similar mass all can strongly influence the Ly forest signal. Therefore, individual lines of sight are dominated by cosmic variance (sourced not only by the large-scale structures of the Universe but also by the ‘structure formation noise’ described above). Current observational efforts are still limited to a handful of sightlines, that however might be not sufficient to extract the intrinsic signal.
@@
In order to test the reliability of current constraints and to guide future observations, we show in Fig. 2 the GaLCC signal extracted from our simulations for different values of the total length of the Ly forest spectra (). In practice, we employ up to 300 spectra of equal length , but decide to report the results using the total spectral length in order to ease the comparison with observations. In each panel, the thin colored lines are the values obtained using different sets of sightlines with total spectral length equal to the value reported in the bottom right part of the panel. For comparison, we show using a thick black line our ‘golden standard’, i.e. the value obtained using all 300 spectra available. It appears immediately clear that this signal is cosmic variance-dominated in the top panels (corresponding to ), while the recovered signal becomes close to the intrinsic one (that we assume is traced by our black thick line) for .
@@
We compare the prediction from our simulation with available observations, namely Meyer et al. (2019, cGpc), Meyer et al. (2020, cGpc), the EIGER JWST program (dashed lines for Kashino et al. 2023 – cMpc –, dot-dashed lines for Kashino et al. 2025 – cMpc) and the ASPIRE JWST program (purple diamonds Kakiichi et al., 2025). Observations from Meyer et al. (2019, 2020) and ASPIRE span a large enough path length to overcome cosmic variance and extract the true reionization signal. It should be noted, however, that to reach such large , they employ somewhat shallow observations, reducing the number of galaxies available, and compound together large redshift intervals, potentially encountering lightcone effects due to the rapid evolution of the ionization field towards the end of the EoR (as discussed in Sec. 3.6).
In the case of Kashino et al. (2023), reporting the results from a single sightline from the EIGER survey, we predict that their results are completely dominated by cosmic variance at , and therefore cannot be quantitatively nor qualitatively trusted as representative of the average GaLCC. The aforementioned results put the findings of Kashino et al. (2023) in a different light. In fact, while they find a qualitative similarity between their measured and the one reported by the thesan simulations at redshift , such agreement is likely entirely driven by cosmic variance. The latest EIGER results (Kashino et al., 2025), instead, include all 6 sightlines and, therefore, correspond to a much larger . This, combined with their deep observations resulting in a rich galaxy sample, makes them much more reliable in a statistical sense. At smaller distances, the flux suppression due to the overdensities harbouring galaxies dominates over cosmic variance and, therefore, the observed signal can be trusted, at least qualitatively. These results are in much better agreement with numerical predictions and present a clear picture of the transition from a galaxy-dominated radiation field to a UVBG-dominated one.
@@
It should be noted that quantitative results depend on the galaxies selected when computing the cross-correlation (or, similarly, detected by a survey). In particular, in Fig. 2 we have selected all galaxies with stellar mass . We provide a visual impression of the impact of such choice in Fig. 3 and 4, where we have selected all galaxies with stellar mass and , respectively. Comparing these figures it appears clear that larger threshold masses increase the cosmic variance in the recovered signal because of the diminishing number of galaxies selected with higher mass thresholds, therefore pushing the requirements to longer to suppress cosmic variance. We have checked that this conclusion is negligibly affected by the choice of binning.
@@ 3.2 What field of view do we need?
Many observational surveys have limited field of view around the sightline. This can negatively impact their ability to study the GaLCC, since it requires us to probe distances up to cMpc in order to capture the extent of the flux enhancement. However, it is important to realise that such distance is not the transverse distance to the sightline (), but rather the 3D distance between a pixel in the spectrum and a galaxy. Therefore, observations can probe 3D separations , at the price of a reduced number of pixel-galaxy pairs. Therefore, here we answer the question: What is the impact of a maximum transverse distance on the cross-correlation signal?
In order to address this question, we show in Fig. 5 the cross-correlation obtained by imposing different values for the maximum galaxy-sightline (transverse) distance (solid colored lines). For comparison, we show the predicted cross-correlation with no maximum limit (black dashed line). It appears clear that a transverse length as small as is sufficient to capture the cross-correlation signal without adding any significant noise. However, limiting severely decreases the number of configurations (i.e. galaxy–pixel pairs) at distances , therefore increasing the statistical noise. Nevertheless, our results are very encouraging for future surveys aiming at detecting this cross-correlation, as even with limited fields of view around the quasar sightline, the signal can be faithfully recovered. In fact, while both probing larger and observing more sightlines increase the number of galaxy–pixel pairs, the latter allows us to probe a larger variety of cosmic environments, and it should therefore be favoured. Finally, limiting also reduces the (average) number of galaxies around each spectrum used in the computation of the GaLCC. We report this number for the same cuts in the outer colorbar. This, in combination with Fig. 2, 3, 4, 9 and 10 (and relative discussions) gives a comprehensive overview of the impact of galaxy selection on the GaLCC.
@@ 3.3 How accurate should the spectra be?
We begin by exploring the impact of noise, assumed to be Gaussian and uncorrelated between pixels. The noise level is characterised by the global signal-to-noise ratio (S/N) of the continuum-normalized spectrum. In practice, we add to the normalised flux in each pixel a Gaussian noise with . We do not impose the resulting flux in each pixel to be positive (but show in Appendix C the implication of doing so). Notice that, since our spectra are produced with spectral resolution of km/s, the value of S/N corresponds to the average signal-to-noise ratio per resolution element. For each value of the S/N, we produce 50 different realizations of the Gaussian noise. Fig. 6 shows the impact of S/N on the cross-correlation (with shaded regions reflecting the S/N). The figure shows that the cross-correlation signal is well recovered for S/N as low as 1 of the normalised flux. However, this results emerges only when showing the full envelope of possible values, while for a single noise realisation the true signal is almost-always completely lost in the noise, as discussed already.
@@
We note that in this specific case, the choice of focusing on the signal at has some impact. In fact, this redshift marks a sweet spot in the competition between a larger intrinsic signal (achieved by moving to earlier times Garaldi et al., 2022) and more numerous galaxies to beat down the statistical noise (found at later times due to structure growth). Therefore, the noise requirements will become more stringent at different times, either because of the growing impact of statistical noise or because the intrinsic signal weakens (i.e. the cross-correlation flattens on to the line).
@@
The next property of the spectra that we investigate is the spectral resolution , where is the largest velocity difference that can be distinguished and is the speed of light. For each of the investigated values of we have rebinned the spectra using a boxcar filter to achieve the desired resolution (the original synthetic spectra have resolution km/s). Fig. 7 shows the resulting cross-correlation for the different values, and clearly shows that a spectral resolution of is sufficient to capture the salient features of the cross-correlation at radii . A somewhat more stringent constraints is found at smaller separations (), where a spectral resolution of is required in order to recover the intrinsic cross-correlation. The reason is simply that at such resolution a pixel corresponds to a physical length of , and therefore smaller scales cannot be captured by the cross-correlation.
@@
The loose resolution requirements are very promising, as it renders much easier to increase the number of observed spectra. Unfortunately, the impact of resolution and noise are not independent. In fact, the resilience to noise comes from the averaging process performed while computing the cross-correlation, which averages out the impact of pixel-level noise. However, this only works as long as the regions of enhanced and suppressed transmitted flux in the spectrum contain a sufficient number of pixels. Therefore, when the spectral resolution decreases the resilience to noise worsens. To demonstrate this and to ease the task of striking a balance between the spectral resolution and S/N when attempting to recover the intrinsic signal, we show in Fig. 8 the impact of noise spectra of different resolutions. In practice, we explore the S/N in Fig. 6 (colored bands) and spectral resolutions in Fig. 7 (panels from top to bottom, is indicated in the bottom right of each one of them). As anticipated, lower are more sensitive to noise. However, even for spectral resolutions as low as (second panel from the top), a S/N of 20 (dark blue band) is sufficient to recover the general GaLCC shape, while for (third panel from the top), S/N=10 is sufficient. We remind the reader that the colored bands show the outer envelope of the GaLCC computed for different noise realizations, therefore providing a worst-case scenario. In reality, most of the observed curves will be significantly closer to the noiseless one.
In a fully-realistic observational setting, however, there are other effects that might alter – at least quantitatively – our conclusions. For instance, the OH airglow is a significant foreground to ground-based optical and near-infrared spectroscopic observations. Removing it might substantially increase the spectral resolution needed, in order to reduce the blending of the OH lines with the background signal.
@@
Our results show that even with moderate noise it is possible to recover the salient features of the cross-correlation signal, shining a promising light on the possibility of significantly enlarging the number of LOS usable for this measurement. Clearly, a major issue in using sightlines for this purpose is the need to identify galaxies around them. Overall, these results, combined with the one provided in Sec. 3.1, outline the observational requirements for the reliable characterization of the GaLCC. In Sec. 5 we will discuss in more details which surveys (available or forthcoming) can be used for this measurement.
@@ 3.4 How to select galaxies?
Observations of the GaLCC employ different techniques to identify galaxies, ranging from C IV absorption in the same spectrum to O III emission around the sightline. Therefore, it is unclear how to properly compare such observations among themselves and with simulations. For this reason, here we have decided to remain agnostic and select galaxies based on their stellar mass. However, it is key to determine whether these different selection techniques are biasing the inferred GaLCC. Additionally, it is foreseeable that in the coming years more and more surveys will identify galaxies through their emission lines. The O III emission line is particularly promising, and is in fact used by a number of ongoing JWST surveys. For this reason, we study here the effect of selecting galaxies based on their star formation rate (SFR), stellar mass () and O III flux. For the latter, we employ the synthetic galaxy SEDs described in Kannan et al. (2022b); Garaldi et al. (2024). Thanks to these, we can restrict the galaxy selection to the brightest O III emitters in the simulation box, defined here based on their equivalent width (EWOIII).
While the limited volume of our simulations hinders a full one-to-one comparison with e.g. the EIGER survey, it still provides precious information on the biases carried by such selection. In Fig. 9 we show the cross-correlation computed selecting the top 0.1%, 1%, 20% and 50% (panels from top to bottom) galaxies when ranked in O III flux, SFR, stellar mass and gas metallicity (as a simple proxy for the strength of their C IV absorption signature in the QSO spectrum). Notice that the galaxy SED were computed only for a subset of ‘well-resolved’ galaxies (as defined in Kannan et al., 2022b), resulting in approximately 10000 galaxies at . Therefore, to enable a proper comparison, we have restricted all selection to the same subset of galaxies, despite the SFR, stellar mass and gas metallicity being available for all simulated galaxies. The different panels show a consistent picture, namely that the selection performed does not affect the resulting cross-correlation. More selective criteria result in a more noisy signals (because of the smaller number of objects resulting in larger noise) and higher peaks (because the sources are more biased and more luminous, as discussed in Sec. 4.1 Garaldi et al., 2022), but there is no discernible difference between selecting galaxies based on their stellar mass, star formation rate or O III flux.
@@
The insensitivity of the GaLCC to the source selection stems from two conspiring reasons. First, in the thesan simulation most galaxies lie on the star-formation main sequence and on the mass-metallicity relation (Garaldi et al., 2024), with only very few deviating from it (Shen et al., 2024). Therefore, the most star forming galaxies are also the most massive ones (in terms of stellar component) and the most metal-rich ones. Second, the large volume covered by the simulations does not allow to resolve the ISM of the simulated galaxies. Therefore, the line emission has to be ‘painted’ on the unresolved ISM using approximate methods (Kannan et al., 2022b). This entails that, for the O III line, there is a built-in correlation between its flux and the star formation of the galaxy. While the complex radiation transport and dust obscuration entering in the SED production can break this correlation, it is the case that the O III flux and the galaxy SFR are tightly related in thesan (Kannan et al., 2022b). These two effects conspire to essentially yield the same galaxy sample, regardless of the selection method. In fact, the overlap between the samples of galaxies resulting from the three selection methods discussed is larger than 85% in all but one case. We caution, however, that this result might change if we were able to resolve the ISM of galaxies in a large-enough volume to compute the GaLCC. Unfortunately this is beyond the reach of current simulations.
@@
The results presented above strengthen the approach taken in this paper and in previous theoretical analysis, i.e. selecting galaxies based on their stellar mass. It also provides precious information for future observational studies.
3.5 Impact of gaps in the galaxy catalog and spectrum